OpenStudy (anonymous):

Solve the equation by using the quadratic formula: 9x^2+18x+4=0. . i keep getting a wrong answer. can anybody show me their steps?

OpenStudy (anonymous):

i got (-18+-6√5)/18

OpenStudy (anonymous):

did you use the formula? maybe you got it right but didn't write in simplest radical form

OpenStudy (anonymous):

well i got radical 180. it became 36 and 5. then it became 6 and 5 right?

OpenStudy (anonymous):

that is right. you just are not done

OpenStudy (anonymous):

does it become -1? in the front? and the 18 on the bottom cancels out?

OpenStudy (anonymous):

\[\frac{-18\pm6\sqrt{5}}{18}=\frac{6(-3\pm\sqrt{5})}{18}=\frac{-3\pm\sqrt{5}}{3}\]

OpenStudy (anonymous):

i thought i couldn't touch the 6 in front of the radical.